OpenStudy (anonymous):

A family bought a new house 10 years ago for $210,000. The house is now worth $302,000. Assuming a steady rate of growth, what was the yearly rate of appreciation?

OpenStudy (anonymous):

they need a rate growth though

OpenStudy (anonymous):

you would use the equation y=a (1+r)^t

OpenStudy (tkhunny):

Okay, then do it that way. Seems good. Let's see your work. Constant RATE of inflation is correct. Constant AMOUNT of inflation is not correct.

OpenStudy (anonymous):

Try starting with \[\frac{302}{210}=(1+R)^{10}\] and solve for R.

OpenStudy (tkhunny):

Although, it would make more sense to start with \(302 = 210(1+R)^{10}\).

OpenStudy (anonymous):

I think I solve it ......) y=a (1+r)^t 302,000=210,000 (1+r)^10 √(10&1.4351)=√(10&(1+r)^10 ) 1.0370=1+r -1 -1 .037=r r= 3.7%

OpenStudy (anonymous):

Very good. That's exactly right!

OpenStudy (anonymous):

Thank you!